OpenStudy (anonymous):

Divide 4 into -1/4 and then put into simplest form. Show work.

OpenStudy (swissgirl):

@completeidiot plllllllllzzzzz can u do the explaining

OpenStudy (anonymous):

dividing by a fraction is the same as multiplying by the reciprocal ex. \[\frac{1}{\frac{1}{4}}=1*\frac{4}{1}\] \[\frac{1}{2}\div \frac{3}{4}=\frac{1}{2}*\frac{4}{3}\] \[4 \div -\frac{1}{4}=???\]

OpenStudy (anonymous):

-16

OpenStudy (anonymous):

4*-4 =-16